OpenStudy (anonymous):

Find the derivative f(θ)= secθ^2 How to solve this?

OpenStudy (inkyvoyd):

use the chain rule

OpenStudy (inkyvoyd):

d/dx f(g(x))=f'(g(x))g'(x)

OpenStudy (anonymous):

f(θ) = y f'(θ) = dy/dθ u = sec(θ) du/dθ = sec(θ)tan(θ) y = u² dy/du = 2u Then, f'(θ) = 2usec(θ)tan(θ) ==> 2(sec²(θ))sec(θ)tan(θ) ==> 2sec³(θ)tan(θ) I hope this helps!

OpenStudy (inkyvoyd):

f(x)=x^2 g(th)=sec th

OpenStudy (anonymous):